The first step in the commercial production of titanium metal is the reaction of rutile (TiO2) with chlorine and graphite:

TiO2(s) + 2Cl2(g) + 2C (s)\rightarrow TiCl4(l) + 2CO (g)

Use the following data to calculate Keq for this reaction at 25° C.

Formula ?H°f (KJ/mol) S° (J/K mol)
TiO2(s) -944.0 50.6
Cl2(g) 0 223.0
C (s) 0 5.7
TiCl4(l) -804.2 252.3
CO (g) -110.5 197.6
Answer #1

TiO2(s) + 2Cl2 (g) + 2C (s) \rightarrow TiCl4(l) + 2CO (g)

\DeltaHo = \DeltaHof(products)-\DeltaHof(reactants)

\DeltaHo =\DeltaHof(TiCl4) + 2 x \DeltaHof (CO) - (\DeltaHof(TiO2)+2 x \DeltaHof(Cl2) + 2 x \DeltaHof(C))

\DeltaHo = -804.2 + 2 x (-110.5) -(-944.0 + 2 x 0 +2 x 0)

\DeltaHo = -1025.2 + 944.0 kJ/mol

\DeltaHo = -81.2 kJ/mol

\DeltaSo = \DeltaSof(products)-\DeltaSof(reactants)

\DeltaSo =\DeltaSof(TiCl4) + 2 x \DeltaSof (CO) - (\DeltaSof(TiO2)+2 x \DeltaSof(Cl2) + 2 x \DeltaSof(C))

\DeltaSo = 252.3 + 2 x (197.6) -(50.6 + 2 x 223.0 +2 x 5.7)

\DeltaSo = 647.5 - 508.6 J/K mol

\DeltaSo = 138.9 J/K mol

\DeltaGo = \DeltaHo - T\DeltaSo

\DeltaGo =  -81.2 kJ/mol - 298 x 138.9 J/K mol

\DeltaGo = -122.59 kJ/mol

\DeltaGo = -RT ln Keq

-122.59 x 103 J/mol = - 8.314 J mol-1 K-1 x 298 K x ln Keq

ln Keq = 122590/2477.57

ln Keq = 49.47

Keq = e49.47

Keq = 3.082 x 1021